OpenStudy (anonymous):

>>> Will Medal and Fan! <<< Subtract. –15 – (–12) A. –27 B. –3 C. 3 D. 27

OpenStudy (anonymous):

Since its - i think A.

OpenStudy (anonymous):

Double negative is a positive I believe.

OpenStudy (anonymous):

Really? I never knew that.

OpenStudy (anonymous):

When there is () you must change the sign to a plus and turn the number positive. Though I believe in this case it's negative 3.

OpenStudy (anonymous):

When there is a negative sign in front on parenthesis change the sign of all of the #'s within the parenthesis. -5-(-1)= -5+1=-4

OpenStudy (anonymous):

Thanks a bunch wish i could fan again :(

OpenStudy (anonymous):

No need! I'm just glad I could help <3 If you have any more questions just ask.

OpenStudy (anonymous):

Ok!